The jackson company has invested in a machine that cost $70000, that has a useful life of seven years, and that has no salvage value at the end of its useful life. The machine is being depreciated using the straight line method, based on its useful life. It has a payback period of four years. What Is the simple rate of return on the machine?
Answer #1

10.7%

70,000 / 4 = 17,500 (yearly cash flow)

70,000 / 7 = 10,000 (depreciation per year)

17,500 - 10,000 = $7,500 (annual increment net oper income)

7,500 / 70,000 = 10.7% (simple rate of return)
Answer #2

Answer: 10.7%

70,000 / 4 = 17,500 (yearly cash flow)

70,000 / 7 = 10,000 (depreciation per year)

17,500 - 10,000 = $7,500 (annual increment net oper income)

7,500 / 70,000 = 10.7% (simple rate of return)